OpenStudy (anonymous):

Find a power series representative of the function and determine it's interval of convergence. \[g(x)=\frac{1}{(1-x)^2}\]

OpenStudy (turingtest):

so do you understand how they get the power series representation of the function in that example? http://tutorial.math.lamar.edu/Classes/CalcII/PowerSeriesandFunctions.aspx

OpenStudy (turingtest):

at which point exactly do you feel lost?

OpenStudy (anonymous):

I believe I do. I compare it to equation 2 \[\sum_{n=0}^\infty x^n=\frac{1}{1-x}\] so in this case it's \[\sum_{n=0}^\infty\] somehow I want to do this, make the equation into a similar form and substitute into the sum :